They say the digit 5 is usually rounded up but can also be rounded down. My question is how would you round the numbers in the equation 9.5+4.7+3.2+7.5=x to the nearest whole number without overestimating or underestimating?
Found 2 solutions by CubeyThePenguin, ikleyn:
Answer by CubeyThePenguin(3113)

To get the closest value, round one of the 5's up and round one of the 5's down like this:

9.5 + 4.7 + 3.2 + 7.5 = 24.9

becomes 10 + 5 + 3 + 7 = 25.

Here, we rounded 9.5 up and round 7.5 down. We can also do the opposite (round 9.5 down and round 7.5 up):

9 + 5 + 3 + 8 = 25.

Answer by ikleyn(52781)

.

First, you INCORRECTLY cite the rule of rounding.


Below is THIS RULE in its classic form

    If the number you are rounding is followed by 5, 6, 7, 8, or 9, round the number up. Example: 38 rounded to the nearest ten is 40. ...

    If the number you are rounding is followed by 0, 1, 2, 3, or 4, round the number down. Example: 33 rounded to the nearest ten is 30.


Second, in your example, you MUST make all calculations FIRST, prior the rounding, and then round at the end.


So, if you do it in accordance with the rules and with the common sense, you will not have any problems.


/\/\/\/\/\/\/\/


Ignore the post by @CubeyThePenguin, since his (or her) writing is a GIBBERISH.